OpenStudy (anonymous):

find dy/dx= 9tan(x/3)

zepdrix (zepdrix):

\[\Large y\quad=\quad 9\tan\left(\frac{1}{3}x\right)\]Mmm so we need to find y'?

zepdrix (zepdrix):

Do you recall the derivative of tan x?

OpenStudy (anonymous):

sec^2 x?

zepdrix (zepdrix):

good good c:

zepdrix (zepdrix):

Ok we'll get a similar result here, we'll just need to remember to apply the chain rule:\[\Large y'\quad=\quad 9\sec^2\left(\frac{1}{3}x\right)\color{royalblue}{\left(\frac{1}{3}x\right)'}\]Chain rule tells us to multiply by the derivative of the inner function. So we still need to take the derivative of the blue part. Understand why we ignore the 9 when we take derivative of tan?

OpenStudy (anonymous):

yes (:

OpenStudy (anonymous):

is it just 1/3?

zepdrix (zepdrix):

The blue part? 1/3x changes to 1/3 when we differentiate? Yup sounds right! :)

zepdrix (zepdrix):

Don't forget to simplify! :O

OpenStudy (anonymous):

so then what happens to the front? :o

zepdrix (zepdrix):

So you got this far?\[\Large y'\quad=\quad 9\sec^2\left(\frac{1}{3}x\right)\left(\frac{1}{3}\right)\]Remember that multiplication is commutative, we can multiply in any order. So we can drag this 1/3 to the front and multiply it with the 9 to simplify.

OpenStudy (anonymous):

y prime= 3 sec^2 (1/3x) >?

OpenStudy (anonymous):

lol minues the >

zepdrix (zepdrix):

Yah looks good! without the bird beak. yay team!\c:/

OpenStudy (anonymous):

Yay! :D Mind helping me on a similar one? (:

zepdrix (zepdrix):

Sure

OpenStudy (anonymous):

would y= sin^2(3x-2) = yprime= 3cos ^2(3x-2) ? ._.

zepdrix (zepdrix):

Woops! Let's make sure we understand what the `outermost` function is.\[\Large y\quad=\quad \sin^2(3x-2)\quad=\quad \left[\sin(3x-2)\right]^2\]

zepdrix (zepdrix):

With trig functions, we write the exponents in a weird location, so it can be a little difficult to identify the outer function sometimes. So clearly here, our outer function is ( )^2, right? So we need to start by applying the power rule.

zepdrix (zepdrix):

So power rule, then chain rule:\[\Large y'\quad=\quad 2\left[\sin(3x-2)\right]\color{royalblue}{\left[\sin(3x-2)\right]'}\]

OpenStudy (anonymous):

so instead the front would be 6? :)

zepdrix (zepdrix):

Yes a 6, but we also get some other business going on :o Does that last step make sense? It's a bit tricky.

OpenStudy (anonymous):

no :o whats with that? lol

zepdrix (zepdrix):

This is an important little tidbit to remember:\[\Large \sin^2x\quad=\quad (\sin x)^2\]^ Do you understand how to take the derivative of this function just above?

OpenStudy (anonymous):

yes just bring it down?

OpenStudy (anonymous):

2 cos x?

zepdrix (zepdrix):

Mmm not quite. We start by bringing the 2 down, as you said. `But we leave the inside alone while we're changing the outside`. \[\Large 2(\sin x)\]But chain rule tells us we need to multiply by the derivative of the inside:\[\Large 2(\sin x)\color{royalblue}{(\sin x)'}\] Are you starting to see the difference? We can't apply the power rule and derivative of sine at the same time. Outer function is ( )^2, we apply power rule. Then we multiply (on the outside) by the derivative of the inner part.

zepdrix (zepdrix):

Uh oh :( I heard that... the sound of your head esploding... ugh

OpenStudy (anonymous):

lol (: no i guess i kind of get it

OpenStudy (anonymous):

kinda

zepdrix (zepdrix):

Umm here's another example that might help.\[\Large (3x+1)^2\]If we take the derivative, it doesn't change to:\[\Large 2(3)\]When we bring the 2 down, we should still have (3x+1) inside, right? We can't apply the power rule AND the 3x+1 at the same time. It leads us down the wrong path.

zepdrix (zepdrix):

Here's the derivative of this example just to compare:\[\Large \left[(3x+1)^2\right]'\quad=\quad 2(3x+1)(3)\]

OpenStudy (anonymous):

oh okay(: i get that one!

OpenStudy (anonymous):

so the equation would be set up as 2 sin(3x-2) x 2cos(3x-2) x 3 ?

OpenStudy (anonymous):

ok my head is exploding lol

zepdrix (zepdrix):

Ahhh you're so close, you have an extra 2 for some reason.\[\Large 2\sin(3x-2)\]Then we multiply by the derivative of the inner function (which in this case is sin(3x-2) which becomes cos(3x-2) ):\[\Large 2\sin(3x-2)\cos(3x-2)\] Then we multiply by the derivative of the inner function again (which in this case is (3x-2) which becomes (3) ): \[\Large 2\sin(3x-2)\cos(3x-2)(3)\]

OpenStudy (anonymous):

ohhhh okay

OpenStudy (anonymous):

got that

OpenStudy (anonymous):

6 sin(3x-2)cos(3x-2) ?

zepdrix (zepdrix):

Yes

OpenStudy (anonymous):

thank you :)